The follwoign data have been obtained for the effect of solvent composition on the solubility of a serine, an amino acid, at 10.0 degrees C:

Volume % Methanol 0 10 20 30 40 60 80 100

Solubility (g/100mL solvent) 22.72 18.98 11.58 6.415 4.205 1.805 0.85 0.65

Solution Density (g/mL) 1.00 0.98 0.97 0.95 0.94 0.91 0.88 0.79

The data were obtained by mixing known volumes of methanol and water to obtain the desired solven compositions and then slowly adding measured amounts of serine to each mixture until no more would go into solution. The temperature was held constant at 10.0

Derive an expression for solvent composition expressed as mass fraction of methanol, x, as a function of volume fraction of methanol, f.
Answer #1

(a) We can derive the relation as

Mass fraction of methanol = xm

Volume fraction of methanol = xv

Mass of methanol = Mm

Volume of methanol = Vm

Mass of water = Mw

Volume of water =Vw

Therefore xm = Mm / Mm + Mw = dm X Vm / dmVm +dwVw

dm= density of methanol

dw = density of water

Also xv = Vm / VT

VT = Total volume

xm = dm * xv *V / dm xv V + dw (1-xv) V